OpenStudy (anonymous):

Examine the function for relative extrema and saddle points. g(x,y)=x^2-y^2-x-y

OpenStudy (anonymous):

\[g(x,y)=x^2-y^2-x-y\] I understand how to do the problem, I just want to make sure I've gotten the right critical points.

OpenStudy (anonymous):

After I got the derivatives setting them equal to zero, I get: \[g_{x}(x,y)=2x-1=0\] \[g_{y}(x,y)=-2y-1=0\]

OpenStudy (anonymous):

factoring out y should be 2(-y-1)=0, so I get y=, so the critical points are (1,1) right? or did I factor this one wrong?